Please help with all the steps shown. I have been agonizing over this one. Thanks in advance:)
Wal-Mart sells thermometers ($2)and hot-water bottles ($6). In December, Wal-Mart's total sakes were $1200. Customers bought 7 times as many thermometers as hot-water bottles. How many of each did Wal-Mart sell?

Found 2 solutions by mmm4444bot, Paul:
Answer by mmm4444bot(95)

Hello There:
This problem asks for two unknown amounts, so we need to come up with expressions to represent them. Since we are told that the relationship between the sales of thermometers and hot-water bottles is that the number of thermometers sold is seven times the number of hot-water bottles, we can assign the following expressions for the number of each.
x = the number of hot-water bottles sold
7*x = the number of thermometers sold
The amount of money that Wal-Mart collects from the sale of each is expressed as PRICE times QUANTITY.
So, the money collected for selling x number of hot-water bottles is 6*x.
The money collected for selling 7*x thermometers is 2*(7*x) which is 14*x
We can now write an equation because we know what the total amount of money collected is:
6*x + 14*x = 1200
Combine like terms on the left side:
20*x = 1200
Solve for x by dividing both sides by 20:
x = 1200/20 = 60
Remember what x stands for. This result tells us that 60 hot-water bottles were sold.
7*x = 420
There were 420 thermometers sold.
Let's check these results.
$2*420 + $6*60 = $840 + $360 = $1,200
It checks out; our answer is correct.
~ Mark

Answer by Paul(988)

Let x be the hot water bottels
2(7x)+6(x)=1200
14x+6x=1200
20x=1200
x=60
7 x 60=420
Therefore, 60 hot water-bottels and 420 thermometers were bought.